OpenStudy (anonymous):

A 20 foot ladder is placed against a building to reach a window that is 16 feet above the ground. How many feet away from the building is the bottom of the ladder? 3 4 12 15

OpenStudy (anonymous):

use pythagorus theorum

OpenStudy (anonymous):

what r the numbers for eac side ?

OpenStudy (anonymous):

\[20^2=16^2+b^2\]

OpenStudy (anonymous):

400/256=b

OpenStudy (anonymous):

that means 15

OpenStudy (anonymous):

i got 12?

OpenStudy (anonymous):

its not correct

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

12 was the right answer

OpenStudy (anonymous):

A trig solution:\[20\text{ Cos}\left[\text{ArcSin}\left[\frac{16}{20}\right]\right]=12 \]